OpenStudy (anonymous):

Express answers in simplest exact form. A square with sides of 3sqrt2 is inscribed in a circle. What is the area of one of the sectors formed by the radii to the vertices of the square? A = PI

OpenStudy (anonymous):

A= ? PI

OpenStudy (anonymous):

???

OpenStudy (anonymous):

since sides of are \[3\sqrt{2}\] Diagonal length of the square is 6 that is nothing but the diameter of the circle therefore required area = (area of the circle - area of the square)/4 = (9 . PI - 18) /4 =2.56